# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

### v6

@@ IV-A Overall structure
To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods. As illustrated in the Fig. 6, the network comprises three main components: a backbone network, a neck network, and a detection head. To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network. In the neck network, we utilize HRNetV2[48] to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
@@
Figure 7: Backbone Network with Global Feature Fusion. The Swin Transformer’s window attention mechanism mainly focuses on local features. Although its layered structure increases the receptive field, it may still fall short in understanding foggy scenes. To address this, we designed a Global Feature Fusion Module (GFFM), replacing the final stage’s window attention with standard multi-head attention to capture global relationships.
IV-B Global Feature Fusion Module
In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision. With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information. As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks[49, 50, 51]. However, applying Transformers to the visual domain still faces some challenges. Firstly, traditional Transformer architectures lack the hierarchical structure and down-sampling operations found in CNN models, making them less adaptable to changes in resolution and receptive field, which may hinder handling of complex visual entities. Additionally, calculating global attention at high resolutions can be computationally intensive.
To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers[52, 53, 54]. Among them, the Swin Transformer[52] combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework. As the Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy, we chose the Swin Transformer as our backbone network. However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding. Therefore, as illustrated in the Fig. 7, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration. Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs. At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
IV-C Kernel Feature Fusion Module
Dynamic convolution adjusts convolutional kernel parameters based on input data, enhancing network flexibility and generalization compared to traditional convolution. CondInst[55] is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency. Similarly, CondLane[39] applies dynamic convolution for lane instance segmentation, achieving strong performance.
@@ IV-C Kernel Feature Fusion Module
IV-D Low-level Edge Enhanced Module
@@
TABLE II: Comparison with Advanced Lane Detection on FoggyLane dataset.
  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN[8] AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD[11] ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD[11] ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT[17] CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT[17] CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT[17] CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR[34] WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA[31] AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA[31] AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane[39] ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane[39] ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane[39] ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet[18] CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet[18] CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet[18] CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet[18] CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet[19] IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet[19] IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet[35] IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet[35] IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane[32] IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.60 70.0 PolarRCNN[21] IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.04 171.6 PolarRCNN[21] IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.05 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
@@
Figure 9: Visualization results of CLRNet[18], FLAMNet[19], PolarRCNN[21] and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
TABLE III: Comparison with Advanced Lane Detection on FoggyCULane dataset
  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN[8] AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD[11] ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD[11] ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT[17] CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT[17] CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT[17] CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR[34] WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA[31] AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA[31] AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane[39] ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane[39] ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane[39] ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet[18] CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet[18] CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet[18] CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet[18] CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet[19] IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet[19] IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet[35] IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet[35] IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane[32] IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN[21] IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN[21] IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.72$\uparrow$)
Figure 10: Visualization results of CLRNet[18], FLAMNet[19], PolarRCNN[21] and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
IV-E Loss functions
@@
IV-E1 Lane start point prediction
@@
$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$
IV-E2 Row-based segmentation prediction
@@
IV-E3 Vertical range prediction
@@ IV-E6 Total loss
V Experiments
V-A Datasets
@@
V-B Evaluation Metric
Lane detection is commonly treated as a semantic segmentation task, evaluated using standard metrics such as True Positive (TP), False Positive (FP), False Negative (FN), and True Negative (TN). However, due to its unique structure, lane detection also adopts specialized metrics. Following SCNN’s[8] protocol, we connect labeled or predicted lane points into line segments and expand them into 30-pixel-wide areas. A prediction is considered correct if the Intersection over Union (IoU) with the ground truth exceeds 0.5. Based on this, we compute Precision, Recall, and F1-score, with the following formulas:
@@ V-B Evaluation Metric
V-C Implementation Details

### v7

@@ 4.1 Overall structure
To address the challenge of lane detection in foggy conditions, where visibility is significantly reduced, we propose a fog-enhanced lane detection network inspired by row-wise methods. As illustrated in the Fig. 8, the network comprises three main components: a backbone network, a neck network, and a detection head. To mitigate the loss of global information in foggy images, we introduce a global feature fusion module within the backbone network. In the neck network, we utilize HRNetV2(Wang et al., 2021) to further integrate multi-scale features from different stages of the backbone, enabling the aggregation of information across various scales and resolutions.
@@
Figure 9: Backbone Network with Global Feature Fusion.
4.2 Global Feature Fusion Module
In recent years, Transformer models have achieved significant success in natural language processing, leading researchers to explore their applications in computer vision. With self-attention mechanisms, Transformer models can capture dependencies between different regions and pixels within an image, thus facilitating a better understanding of global image information. As a result, Transformer models have demonstrated outstanding modeling capabilities and performance improvements in various traffic vision tasks(Zhang et al., 2025a; Chen et al., 2023; Zhang et al., 2024b). However, applying Transformers to the visual domain still faces some challenges. Firstly, traditional Transformer architectures lack the hierarchical structure and down-sampling operations found in CNN models, making them less adaptable to changes in resolution and receptive field, which may hinder handling of complex visual entities. Additionally, calculating global attention at high resolutions can be computationally intensive.
To address these issues, researchers have proposed several enhanced architectures, such as hierarchical Transformers(Liu et al., 2021c, 2022; Dong et al., 2022). Among them, the Swin Transformer(Liu et al., 2021c) combines CNN-inspired concepts, integrating local perception with global attention in a hierarchical Transformer framework. As the Swin Transformer has shown strong performance in classification, detection, and segmentation tasks, balancing computational efficiency with accuracy, we chose the Swin Transformer as our backbone network. However, since the Swin Transformer’s window attention mechanism primarily focuses on local features, its hierarchical structure may still lack sufficient global context for foggy scene understanding. Therefore, as illustrated in the Fig. 9, we designed a Global Feature Fusion Module (GFFM) to enhance the network’s global information integration. Specifically, we replaced the window attention mechanism in the final stage with a standard multi-head attention mechanism to capture global correlations across all inputs. At this stage, image resolution has been sufficiently reduced, making the computational cost of global attention manageable.
4.3 Kernel Feature Fusion Module
Dynamic convolution adjusts convolutional kernel parameters based on input data, enhancing network flexibility and generalization compared to traditional convolution. CondInst(Tian et al., 2023) is an instance segmentation framework using dynamic convolution to generate kernel parameters for each instance, reducing network load and improving efficiency. Similarly, CondLane(Liu et al., 2021a) applies dynamic convolution for lane instance segmentation, achieving strong performance.
@@ 4.3 Kernel Feature Fusion Module
4.4 Low-level Edge Enhanced Module
@@
Table 2: Comparison with Advanced Lane Detection on FoggyLane dataset.
  Method Venue Backbone F1@50 F1@65 F1@75 F1@85 mF1 Gflops(G) FPS SCNN(Pan et al., 2018) AAAI 2018 VGG16 84.01 62.08 56.34 12.76 41.78 328.4 18.7 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 81.76 56.54 32.32 6.50 38.24 8.4 327.3 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 82.75 61.25 34.44 7.89 40.23 16.9 176.9 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 81.98 59.61 34.78 7.69 40.06 9.3 247.7 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 84.32 63.08 38.91 10.76 42.63 18.0 167.1 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 86.25 72.42 51.50 17.26 48.95 70.5 28.6 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 83.79 68.23 58.29 25.32 48.91 - 124.4 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 85.19 65.89 61.89 14.57 48.75 41.0 76.7 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 86.24 69.22 65.27 16.55 52.34 43.0 53.0 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 90.38 75.08 47.50 47.50 10.96 10.2 194.7 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 90.58 78.45 56.48 15.59 52.25 19.6 136.5 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 91.54 82.09 65.30 30.68 57.76 44.8 60.2 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 90.96 83.23 68.53 34.78 59.81 11.9 105.0 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 91.56 82.43 65.61 30.20 58.14 21.5 90.0 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 91.58 85.19 73.38 41.02 64.47 42.9 47.7 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 92.09 83.98 72.15 41.61 62.28 18.5 81.3 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 89.94 78.90 55.32 12.96 50.99 30.1 32.0 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 91.16 82.77 64.95 24.22 56.70 21.7 31.7 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 89.61 81.27 65.27 26.15 56.52 22.64 147.6 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 90.02 83.44 71.66 40.51 60.94 32.1 129.0 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 91.99 84.23 68.08 35.27 34.38 18.60 70.0 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 91.07 79.38 57.44 14.74 52.42 19.04 171.6 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 90.02 83.44 71.66 40.51 51.97 16.05 150.4 Ours SwinGFFM-t 92.34 84.65 70.99 35.77 60.82 12.0 70.4 / 304.7* Ours - SwinGFFM-s 95.04(2.95$\uparrow$) 88.61(3.42$\uparrow$) 77.32(3.94$\uparrow$) 45.02(3.41$\uparrow$) 65.60(3.13$\uparrow$) 20.3 38.2 / 192.9*
@@
Figure 11: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyLane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
4.5 Loss functions
@@
4.5.1 Lane start point prediction
@@
$\mathcal{L}_{\text{hm}}=-\frac{1}{N}\sum_{x,y}\begin{cases}(1-P_{xy})^{\alpha}\log(P_{xy}),&\text{if }GT(x,y)=1\\ (1-GT(x,y))^{\beta}(P_{xy})^{\alpha}\log(1-P_{xy}),&\text{otherwise}\end{cases}$ (5)
4.5.2 Row-based segmentation prediction
@@
4.5.3 Vertical range prediction
@@ 4.5.6 Total loss
Table 3: Comparison with Advanced Lane Detection on FoggyCULane dataset
  Method Venue Backbone Normal Crowded Dazzle light Shadow No line Arrow Curve Cross Night Total SCNN(Pan et al., 2018) AAAI 2018 VGG16 91.07 71.32 67.23 65.23 49.14 88.06 67.07 1344 70.78 72.94 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 87.57 65.47 55.21 60.38 38.66 82.26 57.29 1849 60.61 67.69 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 89.10 68.29 60.91 62.03 41.98 85.17 63.88 2371 70.13 70.13 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 90.52 71.52 65.70 68.92 47.31 86.32 62.34 870 67.64 74.30 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 91.68 72.93 67.23 65.23 49.14 88.06 67.07 1344 70.78 75.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 91.12 74.68 68.87 78.61 47.33 85.76 63.88 1008 69.72 76.16 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 85.79 63.45 56.73 58.72 38.20 79.31 55.34 1068 57.62 65.64 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 90.97 72.32 67.31 74.01 46.52 86.32 65.26 1536 73.62 74.97 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 92.09 73.49 67.40 73.87 47.92 87.48 69.28 1722 69.85 75.19 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 91.56 74.41 71.12 77.89 48.59 86.73 71.34 1262 71.00 76.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 91.76 75.55 69.62 74.62 50.01 87.40 71.29 1427 72.11 77.06 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 92.47 76.26 70.91 78.50 51.84 88.93 72.51 1597 72.75 77.90 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 92.55 76.47 71.89 74.45 50.82 88.94 65.93 1122 72.86 78.04 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 92.96 76.32 73.40 79.06 51.46 89.39 71.83 1301 73.64 78.36 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 92.95 76.67 71.49 79.89 52.47 89.29 68.72 1101 74.03 78.72 CLRNet(Zheng et al., 2022) CVPR 2022 DLA34 93.23 77.64 73.90 79.23 52.11 89.45 72.48 1109 73.63 79.04 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet34 92.60 76.61 71.42 79.81 50.89 89.04 71.47 1073 74.17 78.45 FLAMNet(Ran et al., 2023) IEEE TITS 2023 DLA34 92.98 77.26 71.84 79.76 51.29 89.34 71.38 1143 73.93 78.77 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 84.91 69.27 60.86 67.94 45.69 79.04 67.50 1045 68.57 71.33 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 85.61 69.35 61.22 66.50 45.73 79.19 66.51 837 69.93 71.87 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 90.85 73.32 67.18 75.32 47.18 87.14 63.38 1456 69.20 74.61 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 93.34 77.76 75.26 79.76 53.25 89.62 75.94 1289 74.54 78.68 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 92.90 77.50 74.03 80.89 51.71 89.76 73.02 1226 74.21 79.13 Ours SwinGFFM-t 92.68 77.50 71.52 80.42 50.95 88.92 72.08 950 73.11 78.66 Ours - SwinGFFM-s 93.40(0.17$\uparrow$) 78.63(1.99$\uparrow$) 73.89 82.74(2.85$\uparrow$) 53.75(1.28$\uparrow$) 89.82(0.37$\uparrow$) 74.09(1.58$\uparrow$) 1146 74.97(0.80$\uparrow$) 79.85(0.72$\uparrow$)
Figure 12: Visualization results of CLRNet(Zheng et al., 2022), FLAMNet(Ran et al., 2023), PolarRCNN(Wang et al., 2025) and our method on FoggyCULane testing set. The arrows indicate instances of lane lines that were not correctly detected by the corresponding method.
5 Experiments
5.1 Datasets
@@
5.2 Evaluation Metric
Lane detection is commonly treated as a semantic segmentation task, evaluated using standard metrics such as True Positive (TP), False Positive (FP), False Negative (FN), and True Negative (TN). However, due to its unique structure, lane detection also adopts specialized metrics. Following SCNN’s(Pan et al., 2018) protocol, we connect labeled or predicted lane points into line segments and expand them into 30-pixel-wide areas. A prediction is considered correct if the Intersection over Union (IoU) with the ground truth exceeds 0.5. Based on this, we compute Precision, Recall, and F1-score, with the following formulas:
@@ 5.2 Evaluation Metric
Table 4: Comparison with Advanced Lane Detection on FoggyTusimple dataset.
    Method Venue Backbone F1(%) ACC(%) FP(%) FN(%) SCNN(Pan et al., 2018) AAAI 2018 VGG16 94.64 95.89 5.87 4.78 UFLD(Qin et al., 2020) ECCV 2020 ResNet18 94.57 93.60 3.49 7.59 UFLD(Qin et al., 2020) ECCV 2020 ResNet34 95.05 95.10 4.84 5.08 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet18 95.77 95.78 3.83 3.53 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet34 96.15 95.50 4.59 3.06 LaneATT(Tabelini et al., 2021a) CVPR 2020 ResNet122 93.24 95.43 9.61 3.50 LSTR(Liu et al., 2021b) WACV 2021 ResNet34 94.70 95.58 5.74 4.81 RESA(Zheng et al., 2021) AAAI 2021 ResNet34 95.22 95.60 4.63 4.95 RESA(Zheng et al., 2021) AAAI 2021 ResNet50 95.51 95.83 4.53 4.43 CondLane(Liu et al., 2021a) ICCV 2021 ResNet18 95.41 94.81 3.80 5.46 CondLane(Liu et al., 2021a) ICCV 2021 ResNet34 95.65 94.69 3.63 5.13 CondLane(Liu et al., 2021a) ICCV 2021 ResNet101 96.03 95.14 2.94 5.09 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet18 96.48 95.64 3.89 3.11 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet34 96.23 95.94 4.34 3.17 CLRNet(Zheng et al., 2022) CVPR 2022 ResNet101 96.13 95.64 4.15 3.58 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet18 96.52 95.65 3.84 3.09 FLAMNet(Ran et al., 2023) IEEE TITS 2023 ResNet101 96.34 95.98 4.45 2.82 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet18 95.71 93.89 3.53 5.12 DBNet(Dai et al., 2024) IEEE TIV 2024 ResNet34 95.61 93.72 3.48 5.38 HWLane(Zhao et al., 2024) IEEE TITS 2024 Res34 95.54 95.91 4.32 4.60 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 ResNet34 96.33 95.91 4.50 2.77 PolarRCNN(Wang et al., 2025) IEEE TITS 2025 DLA34 96.37 94.99 3.21 4.09 Ours SwinGFFM-t 96.63 95.86 2.77 4.04 Ours - SwinGFFM-s 96.95(0.43$\uparrow$) 96.10(0.12$\uparrow$) 2.32(0.62$\downarrow$) 3.83
5.3 Implementation Details